# Circle bundles with PSC over large manifolds

### v3

@@ 3.2. Proof of Theorem B
    If the circle bundle $E$ is not fiberwise orientable, then $w_{1}(E)\neq 0\in H^{1}(M;\mathbf{Z}_{2})$. By using the isomorphism $H^{1}(M;\mathbf{Z}_{2})\cong\mathrm{Hom}(\pi_{1}(M),\mathbf{Z}_{2}),$ we obtain a homomorphism $\phi:\pi_{1}(M)\to\mathbf{Z}_{2}$ corresponding to $w_{1}(E)$. We pass to the double cover of $M$ corresponding to the index $2$ subgroup $\ker(\phi)$ of $\pi_{1}(M)$. The pullback of $E$ to this cover is an oriented circle bundle by construction. By the reasoning in Part $(1)$, the total space of the pulled back bundle continues to be PSC. If $\dim_{mc}(\widetilde{M})=3$, then the the macroscopic dimension of this cover is also $3$. If $M$ is enlargeable, then the cover is also enlargeable: it admits a degree $2$ map to $M$ and it is spin since all orientable $3$-manifolds are spin. Hence, the cover is also enlargeable. Therefore, we may also assume without loss of generality that the bundle $E\to M$ is fiberwise orientable.
By Proposition 8, $M$ contains an aspherical prime summand $K$, so that $M=K\#N$ for some $3$-manifold $N$. Recall that fiberwise oriented circle bundles $E\to M$ are classified by the Euler class $e(E)\in H^{2}(M;\mathbf{Z})$. Note that
$H^{2}(M;\mathbf{Z})\cong H^{2}(K;\mathbf{Z})\oplus H^{2}(N;\mathbf{Z})$
Therefore, $E$ is isomorphic to a fiber connect sum of two circle bundles over $K$ and $N$, respectively. Let $L\to K$ denote the circle bundle corresponding to the aspherical summand $K$. Therefore, the total space of $E$ admits a degree $1$ map to the total space of $L$. Since the total space of a circle bundle over an aspherical manifold is aspherical, this implies that the total space of $E$ admits a degree $1$ map to an aspherical $4$-manifold. Hence, the total space of $E$ can not admit a positive scalar curvature metric by a result of Chodosh, Li and Liokumovich [CLL23, Corollary 3]. This concludes the proof. ∎
@@ 4. Further examples and remarks
We remark that using Example 9, we can produce examples of macroscopically large manifolds $Z$ admitting circle bundles $p:E\to Z$ such that the total space $E$ is macroscopically small, yet the asymptotic dimension of $\pi_{1}(E)$ is large. This indicates that the macroscopic smallness of $E$ cannot be deduced from just the fundamental group. [^4] For any metric space $X$, $\dim_{mc}(X)\leq\mathrm{asdim}(X)$. For more information on the asymptotic dimension, we refer the reader to the survey [BD08].
@@
Let $m\geq 2$, and $Z^{2m}\subset(T^{2n},\omega)$ be a symplectic submanifold obtained by repeated applications of Theorem 1. Hence, by Proposition 2, $\pi_{1}(Z)\cong\mathbf{Z}^{2n}$. By Part $(2)$ of Theorem A, there exists a circle bundle $S^{1}\hookrightarrow E\stackrel{{\scriptstyle p}}{{\to}}Z$ such that $\dim_{mc}(\widetilde{E})\leq n+1$. From the homotopy long exact sequence for fiber bundles, we obtain an exact sequence
$0\cong\pi_{2}(E)\to\pi_{1}(S^{1})\cong\mathbf{Z}\to\pi_{1}(E)\to\pi_{1}(Z)\to 0$
Therefore, $\pi_{1}(E)$ is an infinite cyclic extension of $\pi_{1}(Z)\cong\mathbf{Z}^{2n}$. Therefore, the asymptotic dimension of $\pi_{1}(E)$ is $\mathrm{asdim}(\pi_{1}(E))=2n+1$. Likewise, for the virtual cohomological dimension, we have $\mathrm{vcd}(\pi_{1}(E))=2n+1$. Note that for $m=n$, we get $\mathrm{vcd}(\pi_{1}(E))=\dim E$. Thus, in this case the techniques of [BD16] are unable to provide apriori upper bounds on $\dim_{mc}(E)$, even if $E$ is totally non-spin.
@@
Let $N^{2n}$ be an even dimensional enlargeable manifold, for $n\geq 3$. By [FP21, Theorem 1], there exists a symplectic manifold $(M^{2n},\omega)$ and a positive degree map $f:M\to N$. In fact, we may arrange $c_{1}(M)$ to be even by ensuring that in the proof of [FP21, Theorem 1], the chain of symplectic submanifolds of the twistor space is selected so that each of them is a Donaldson divisor that is Poincaré dual to an even (respectively, odd) multiple of the symplectic form in the previous submanifold, depending on if the parity of $n$ is even (respectively, odd). This ensures $c_{1}(M)$ is even by the remark at the end of the article [FP21, pg. 3] regarding the first Chern class of the twistor space. Therefore, we have
@@
Let $(M^{2n},\omega)$ be an integral symplectic manifold which is symplectically aspherical such that $\pi_{1}(M)$ is amenable. Let $\omega_{0}$ denote the standard area form on $S^{2}$. Then $(M\times S^{2},\omega\oplus\omega_{0})$ is not symplectically aspherical, hence it fails the hypothesis of Part $(3)$ of Theorem A.
Nevertheless, let $Z^{2n}\subset(M\times S^{2},\omega\oplus\omega_{0})$ be a Donaldson divisor. Consider the map $f:M\times S^{2}\to M$ given by the projection to the first factor. By Proposition 5, $f|_{Z}:Z\to M$ is a positive degree map. Since $M$ is symplectically aspherical, it is rationally essential due to Proposition 7. Since $f|_{Z}$ has positive degree, $Z$ must also be rationally essential. By Proposition 2, $\pi_{1}(Z)\cong\pi_{1}(M)$. Therefore, $Z$ is a rationally essential manifold with amenable fundamental group. By [Dra11, Theorem 7.2], $\dim_{mc}(\widetilde{Z})=2n$.
@@
Let $M=T^{2n}\#\overline{\mathbf{CP}}^{n}$ for $n\geq 3$. By the symplectic blowup construction due to Gompf [Gom95], $M$ admits a symplectic structure $\omega$. Let $\overline{\mathbf{CP}}^{1}$ denote the exceptional divisor. Then, $\overline{\mathbf{CP}}^{1}\subset(M,\omega)$ is a symplectic submanifold. Therefore,
@@
Nevertheless, let $Z^{2n-2}\subset(M,\omega)$ be a Donaldson divisor with $[Z]=\mathrm{PD}(k[\omega])$, for some $k\gg 1$. Consider the map $f:M\to T^{2n-2}$ given by composing the blow-down map $q:M\to M/\overline{\mathbf{CP}}^{1}\cong T^{2n}$ with the projection $T^{2n}\cong T^{2n-2}\times T^{2}\to T^{2n-2}$. For $z\in T^{2n-2}$,
@@
Choose $z\in T^{2n-2}$ such that $\{z\}\times T^{2}\subset T^{2n}$ is disjoint from a neighborhood of the blown up point $\{z_{0}\}=q(\overline{\mathbf{CP}}^{n})\subset T^{2n}$. By the construction of symplectic blow-ups [Gom95], $q$ restricts to a symplectomorphism on the complement of a neighborhood of the exceptional divisor $\overline{\mathbf{CP}}^{1}\subset M$. Therefore, $T^{2}\cong q^{-1}(\{z\}\times T^{2})\subset(M,\omega)$ is a symplectic submanifold intersecting $Z^{2n-2}$ transversely. Therefore, the oriented intersection number is given by:
#$\#(q^{-1}(\{z\}\times T^{2})\cap Z)=k\int_{q^{-1}(\{z\}\times T^{2})}\omega>0$
Thus, $f|_{Z}:Z\to T^{2n-2}$ is a map of positive degree. Hence, $Z$ is rationally essential. By Proposition 2, $\pi_{1}(Z)\cong\pi_{1}(M)\cong\mathbf{Z}^{2n}$ which is amenable. Therefore, by [Dra11, Theorem 7.2], $\dim_{mc}(\widetilde{Z})=2n-2$. Additionally, if $n=3$ (resp. $n=4$), then $Z$ admits a positive degree map to $T^{4}$ (resp. $T^{6}$). In that case, $Z$ is not PSC. [SY79]
@@
Let $(M^{4},\omega)$ be an integral symplectic $4$-manifold with $b_{2}^{+}(M)>1$, and let $\omega_{0}$ denote an integral area form on the surface $S_{g}$. Let $Z^{4}\subset(M\times S_{g},\omega\oplus\omega_{0})$ be a Donaldson divisor. By Proposition 5, the projection $f:Z\to M$ has positive degree. Therefore, the induced map
@@
We close this section with an example which illustrates that macroscopic dimensions of manifolds having diffeomorphic universal covers need not be the same. For more surprising examples, we refer the reader to Kastenholz-Reinhold [KR23].
@@
Let $M=T^{2n}$ for $n\geq 3$ and $\omega$ be the standard symplectic form on the torus, normalized to be integral. Let $Z\subset(M,\omega)$ be a Donaldson divisor with $[Z]=\mathrm{PD}(k[\omega])$. By Part $(2)$ of Theorem A, there exists a circle bundle $S^{1}\hookrightarrow E\stackrel{{\scriptstyle p}}{{\to}}Z$ such that $\dim_{mc}(\widetilde{E})\leq n$. Let $\widehat{E}\to\widetilde{Z}$ denote the circle bundle obtained by pulling back the bundle $p:E\to Z$ along the universal cover $q:\widetilde{Z}\to Z$. The resulting bundle is classified by an element in
@@
More precisely, it is classified by the cohomology class of the $2$-form $q^{*}\omega$. However, since $(Z,\omega)$ is symplectically aspherical by Example 6, for any map $f:S^{2}\to\widetilde{Z}$,
@@
Therefore, $[q^{*}\omega]=0\in H^{2}(\widetilde{Z};\mathbf{Z})\cong\mathrm{Hom}(\pi_{2}(Z),\mathbf{Z})$. This implies that the bundle $\widehat{E}\to\widetilde{Z}$ is trivial, i.e., $\widehat{E}\cong\widetilde{Z}\times S^{1}$. Since the total space of $\widehat{E}$ is a cover of the total space of $E$, we obtain that $\widetilde{Z}\times S^{1}$ covers the total space of $E$. In particular, $\widetilde{E}$ is diffeomorphic to $\widetilde{Z}\times\mathbf{R}$. Note that $\dim_{mc}(\widetilde{Z})=2n-2$, by Part $(3)$ of Theorem A. Thus, $E$ and $Z\times S^{1}$ have diffeomorphic universal covers, but nevertheless have different macroscopic dimensions.

### v5

@@ 3.2. Proof of Theorem B
    If the circle bundle $E$ is not fiberwise orientable, then $w_{1}(E)\neq 0\in H^{1}(M;\mathbf{Z}_{2})$. By using the isomorphism $H^{1}(M;\mathbf{Z}_{2})\cong\mathrm{Hom}(\pi_{1}(M),\mathbf{Z}_{2}),$ we obtain a homomorphism $\phi:\pi_{1}(M)\to\mathbf{Z}_{2}$ corresponding to $w_{1}(E)$. We pass to the double cover of $M$ corresponding to the index $2$ subgroup $\ker(\phi)$ of $\pi_{1}(M)$. The pullback of $E$ to this cover is an oriented circle bundle by construction. By the reasoning in Part $(1)$, the total space of the pulled back bundle continues to be PSC. If $\dim_{\mathrm{MC}}(\widetilde{M})=3$, then the the macroscopic dimension of this cover is also $3$. If $M$ is enlargeable, then the cover is also enlargeable: it admits a degree $2$ map to $M$ and it is spin since all orientable $3$-manifolds are spin. Hence, the cover is also enlargeable. Therefore, we may also assume without loss of generality that the bundle $E\to M$ is fiberwise orientable.
Let $P_{1},\cdots,P_{\ell}$ be all the $3$-manifolds with finite fundamental group appearing in the Kneser-Milnor prime decomposition of $M$. By the Kurosh subgroup theorem, the kernel of the homomorphism $\varphi:\pi_{1}(M)\to\pi_{1}(P_{1})*\cdots*\pi_{1}(P_{\ell})\to\pi_{1}(P_{1})\times\cdots\times\pi_{1}(P_{\ell})$ is torsion-free. Replacing $M$ with the finite cover corresponding to $\ker(\varphi)$ if necessary, we may assume without loss of generality that $\pi_{1}(M)$ is torsion-free. By Proposition 8, $M$ must contain at least one aspherical summand. Thus, M=K#NM=K\#N where $K$ is an orientable aspherical $3$-manifold. If the inclusion of the fiber $S^{1}\hookrightarrow E$ is homotopically nontrivial, it follows from [He25, Theorem 1.9] that $E$ is not PSC. Therefore, let us assume the fibers of $E\to M$ are $\pi_{1}$-null in the total space of $E$. Thus, $\pi_{1}(E)\cong\pi_{1}(M)$ is a torsion-free $3$-manifold group. As such, $\pi_{1}(E)$ has cohomological dimension at most $3$. Furthermore, $\pi_{1}(E)$ satisfies the Strong Novikov Conjecture [MOOP08, Theorem 1.1] [^6] This theorem proves that the Baum-Connes Conjecture with coefficients holds for an oriented $3$-manifold group $G$. The discussion on [MOOP08, pg. 5] shows that the Strong Novikov Conjecture for $G$ follows if the Baum-Connes Conjecture with coefficients holds for $G$.. Next, observe that as $M$ is a closed orientable $3$-manifold, it is spin. Choosing a connection on $p:E\to M$, we obtain a decomposition $TE\cong V\oplus H$ into the vertical and horizontal subbundles. Note that $H\cong p^{*}TM$ and $V$ is the trivial line bundle over $E$. Therefore, $w_{2}(TE)=w_{2}(H\oplus V)=w_{2}(H)=p^{*}w_{2}(TM)=0$. As $M$ is orientable and $p:E\to M$ is fiberwise orientable, the total space $E$ is also orientable. Since $w_{2}(TE)=0$, this implies the total space $E$ is spin. It now follows from [Bol09, Main Theorem, Remark 1.7] that if $E$ is PSC, the classifying map $f:E\to K(\pi_{1}(E),1)$ is homotopic into the $2$-skeleton of $K(\pi_{1}(E),1)$. However, this contradicts Lemma 3.2. Thus, $E$ is not PSC. This concludes the proof. ∎
@@ 4. Further examples and remarks
We remark that using Example 9, we can produce examples of macroscopically large manifolds $Z$ admitting circle bundles $p:E\to Z$ such that the total space $E$ is macroscopically small, yet the asymptotic dimension of $\pi_{1}(E)$ is large. This indicates that the macroscopic smallness of $E$ cannot be deduced from just the fundamental group. [^7] For any metric space $X$, $\dim_{\mathrm{MC}}(X)\leq\mathrm{asdim}(X)$. For more information on the asymptotic dimension, we refer the reader to the survey [BD08].
@@
Let $m\geq 2$, and $Z^{2m}\subset(T^{2n},\omega)$ be a symplectic submanifold obtained by repeated applications of Theorem 1. Hence, by Proposition 2, $\pi_{1}(Z)\cong\mathbf{Z}^{2n}$. By Part $(2)$ of Theorem A, there exists a circle bundle $S^{1}\hookrightarrow E\stackrel{{\scriptstyle p}}{{\to}}Z$ such that $\dim_{\mathrm{MC}}(\widetilde{E})\leq m+1$}. From the homotopy long exact sequence for fiber bundles, we obtain an exact sequence
${\pi_{2}(Z)}\to\pi_{1}(S^{1})\cong\mathbf{Z}\to\pi_{1}(E)\to\pi_{1}(Z)\to 0$
If $m\geq 3$, then $\pi_{2}(Z)\cong\pi_{2}(T^{2n})\cong 0$ by Proposition 2. Therefore, $\pi_{1}(E)$ is an infinite cyclic extension of $\mathbf{Z}^{2n}$. Therefore, the asymptotic dimension of $\pi_{1}(E)$ is $\mathrm{asdim}(\pi_{1}(E))=2n+1$ by [BD08, Theorem 66], since $\pi_{1}(E)$ is polycyclic with Hirsch length $2n+1$. Likewise, for the virtual cohomological dimension, we have $\mathrm{vcd}(\pi_{1}(E))=2n+1$. If $m=2$, then possibly $\pi_{2}(Z)\to\pi_{1}(S^{1})\cong\mathbf{Z}$ has non-zero image. In that case, $\pi_{1}(E)$ is a finite cyclic extension of $\mathbf{Z}^{2n}$. Then $\mathrm{asdim}(\pi_{1}(E))=2n$ by [BD08, Corollary 54] and since $\mathbf{Z}^{2n}$ is a finite-index subgroup of $\pi_{1}(E)$, $\mathrm{vcd}(\pi_{1}(E))=\mathrm{vcd}(\mathbf{Z}^{2n})=2n$.
@@
Let $N^{2n}$ be an even dimensional closed enlargeable manifold, for $n\geq 3$. By [FP21, Theorem 1], there exists a symplectic manifold $(M^{2n},\omega)$ and a positive degree map $f:M\to N$. We may arrange $c_{1}(M)$ to be an even multiple of $\omega]$ by ensuring that in the proof of [FP21, Theorem 1], the chain of symplectic submanifolds of the twistor space of $N$, is selected so that all (resp. all but one) of them is a Donaldson divisor that is Poincaré dual to an even multiple of the symplectic form in the previous submanifold, depending on whether the parity of $n$ is even (respectively, odd). This ensures $c_{1}(M)$ is even by the remark at the end of the article [FP21, pg. 3] regarding the first Chern class of the twistor space of $N$, together with the expression (2.3) in Proposition 4 for the first Chern class of a Donaldson divisor of a symplectic manifold. Therefore, we have
@@
Let $n\geq 2$. Let $(M^{2n},\omega)$ be a closed integral symplectic manifold which is symplectically aspherical such that $\pi_{1}(M)$ is amenable. Let $\omega_{0}$ denote the standard area form on $S^{2}$. Then $(M\times S^{2},\omega\oplus\omega_{0})$ is not symplectically aspherical, hence it fails the hypothesis of Part $(3)$ of Theorem A.
Nevertheless, let $Z^{2n}\subset(M\times S^{2},\omega\oplus\omega_{0})$ be a Donaldson divisor. Consider the map $f:M\times S^{2}\to M$ given by the projection to the first factor. By Proposition 5, $f|_{Z}:Z\to M$ is a positive degree map. Since $M$ is symplectically aspherical, it is rationally essential due to Proposition 7. Since $f|_{Z}$ has positive degree, $Z$ must also be rationally essential. By Proposition 2, $\pi_{1}(Z)\cong\pi_{1}(M)$. Therefore, $Z$ is a rationally essential manifold with amenable fundamental group. By [Dra11, Theorem 1.6], $\dim_{\mathrm{MC}}(\widetilde{Z})=2n$.
@@
Let $M=T^{2n}\#\overline{\mathbf{CP}}^{n}$ for $n\geq 3$. By the symplectic blowup construction [MS12, pg 324-325], $M$ admits a symplectic structure $\omega$. Let $\overline{\mathbf{CP}}^{n-1}\subset M$} denote the exceptional divisor. Then, $\overline{\mathbf{CP}}^{1}\subset{\overline{\mathbf{CP}}^{n-1}}\subset(M,\omega)$ is a symplectic submanifold. Therefore,
@@
Nevertheless, let $Z^{2n-2}\subset(M,\omega)$ be a Donaldson divisor with $Z]=\mathrm{PD}(k[\omega])$, for some $k\gg 1$. Consider the map $f:M\to T^{2n-2}$ given by composing the blow-down map $q:M\to M/{\overline{\mathbf{CP}}^{n-1}}\cong T^{2n}$ with the projection $T^{2n}\cong T^{2n-2}\times T^{2}\to T^{2n-2}$. For $z\in T^{2n-2}$,
@@
Choose $z\in T^{2n-2}$ such that $\{z\}\times T^{2}\subset T^{2n}$ is disjoint from a neighborhood of the blown up point $\{z_{0}\}=q(\overline{\mathbf{CP}}^{n})\subset T^{2n}$. By the construction of symplectic blow-ups, $q$ restricts to a symplectomorphism on the complement of a neighborhood of the exceptional divisor ${\overline{\mathbf{CP}}^{n-1}}\subset M$. Therefore, $T^{2}\cong q^{-1}(\{z\}\times T^{2})\subset(M,\omega)$ is a symplectic submanifold intersecting $Z^{2n-2}$ transversely. Therefore, the oriented intersection number is given by:
$(q^{-1}(\{z\}\times T^{2})\cap Z)=k\int_{q^{-1}(\{z\}\times T^{2})}\omega>0$
Thus, $f|_{Z}:Z\to T^{2n-2}$ is a map of positive degree. Hence, $Z$ is rationally essential. By Proposition 2, $\pi_{1}(Z)\cong\pi_{1}(M)\cong\mathbf{Z}^{2n}$ which is amenable. Therefore, by [Dra11, Theorem 1.6], $\dim_{\mathrm{MC}}(\widetilde{Z})=2n-2$. Additionally, if $n=3$ (resp. $n=4$), then $Z$ admits a positive degree map to $T^{4}$ (resp. $T^{6}$). In that case, $Z$ is not PSC [SY79].
@@
Let $(M^{4},\omega)$ be a closed integral symplectic $4$-manifold with $b_{2}^{+}(M)>1$, and let $\omega_{0}$ denote an integral area form on the surface $S_{g}$. Let $Z^{4}\subset(M\times S_{g},\omega\oplus\omega_{0})$ be a Donaldson divisor. By Proposition 5, the projection $f:Z\to M$ has positive degree. Therefore, the induced map
@@
We give an example which illustrates that macroscopic dimensions of manifolds having diffeomorphic universal covers need not be the same. For more surprising examples, we refer the reader to Kastenholz-Reinhold [KR23].
@@
Let $M=T^{2n}$ for $n\geq 3$ and $\omega$ be the standard symplectic form on the torus, normalized to be integral. Let $Z\subset(M,\omega)$ be a Donaldson divisor with $Z]=\mathrm{PD}(k[\omega])$. By Part $(2)$ of Theorem A, there exists a circle bundle $S^{1}\hookrightarrow E\stackrel{{\scriptstyle p}}{{\to}}Z$ such that $\dim_{\mathrm{MC}}(\widetilde{E})\leq n$. Let $\widehat{E}\to\widetilde{Z}$ denote the circle bundle obtained by pulling back the bundle $p:E\to Z$ along the universal cover $q:\widetilde{Z}\to Z$. The resulting bundle is classified by an element in
@@
More precisely, it is classified by the cohomology class of the $2$-form $k\cdot q^{*}\omega$}. However, since $(Z,\omega)$ is symplectically aspherical by Example 6, for any map $f:S^{2}\to\widetilde{Z}$,
@@
Therefore, $q^{*}\omega]=0\in H^{2}(\widetilde{Z};\mathbf{Z})\cong\mathrm{Hom}(\pi_{2}(Z),\mathbf{Z})$. This implies that the bundle $\widehat{E}\to\widetilde{Z}$ is trivial, i.e., $\widehat{E}\cong\widetilde{Z}\times S^{1}$. Since the total space of $\widehat{E}$ is a cover of the total space of $E$, we obtain that $\widetilde{Z}\times S^{1}$ covers the total space of $E$. In particular, $\widetilde{E}$ is diffeomorphic to $\widetilde{Z}\times\mathbf{R}$. Note that $\dim_{\mathrm{MC}}(\widetilde{Z})=2n-2$, by Part $(3)$ of Theorem A. Thus, $E$ and $Z\times S^{1}$ have diffeomorphic universal covers, but nevertheless have different macroscopic dimensions.
[…]
Example 16 (md-large divisors).
Let $(T^{2n},\omega)$, $n\geq 2$} be a symplectic torus as in Example 9. Consider the product $(T^{2n}\times S^{2},\omega\oplus\omega_{0})$, as in Example 12. Let $Z^{2n}\subset(T^{2n}\times S^{2},\omega\oplus\omega_{0})$ be a Donaldson divisor. Then there is a positive degree map $f:Z\to T^{2n}$ which induces an isomorphism on $\pi_{1}$, by Example 12. By lifting $f$ to the universal cover, we obtain a map with positive local degree
$\widetilde{f}:\widetilde{Z}\to\widetilde{T^{2n}}\cong\mathbf{R}^{2n}$
Therefore, $\widetilde{f}$ induces a non-zero homomorphism between the locally finite homology groups $\widetilde{f}_{*}:H_{2n}^{lf}(\widetilde{Z};\mathbf{Z})\to H_{2n}^{lf}(\mathbf{R}^{2n};\mathbf{Z})$. Consequently, $\widetilde{f}_{*}[\widetilde{Z}]\neq 0$. Since $\pi_{1}(Z)\cong\mathbf{Z}^{2n}$, we have $T^{2n}=K(\pi_{1}(Z),1)$. Therefore, by [Dra13, Theorem 5.4], $\dim_{\mathrm{mc}}(\widetilde{Z})=2n$. By Parts $(1)$ and $(2)$ of Theorem A, we obtain circle bundles $E\to Z$ such that $\dim_{\mathrm{mc}}(\widetilde{E})\leq\dim_{\mathrm{MC}}(\widetilde{E})\leq n+1$ and $E$ is PSC.